# Amplitude Expansion Phase Field Crystal (APFC) Modeling based Efficient Dislocation Simulations using Fourier Pseudospectral Method

### v2

@@ B.3 Proofs of the Remarks in Section 2
Proof of Remark 2.1.
$\Big{\lvert}\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}\eta(\bm{x})e^{\imath\bm{k_{j}}\cdot\frac{\bm{x}}{\varepsilon}}\,{\rm d}\bm{x}-\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}I_{\eta}e^{\imath\bm{k_{j}}\cdot\frac{\bm{x}}{\varepsilon}}\,{\rm d}\bm{x}\Big{\rvert}$ $=\Big{\lvert}\int_{C_{1}}(\eta(\varepsilon\bm{y})-I_{\eta})e^{\imath\bm{k_{j}}\cdot\bm{y}}\,{\rm d}\bm{y}\Big{\rvert}$
$\leq\Big{\lvert}\Big{(}\int_{C_{1}}\lvert\eta(\varepsilon\bm{y})-I_{\eta}\lvert^{2}\,{\rm d}\bm{y}\Big{)}^{\frac{1}{2}}\Big{(}\int_{C_{1}}\lvert e^{\imath\bm{k_{j}}\cdot\bm{y}}\lvert^{2}\,{\rm d}\bm{y}\Big{)}^{\frac{1}{2}}\Big{\rvert}$
$=\Big{\lvert}\Big{(}\int_{C_{1}}\big{\lvert}\eta(\varepsilon\bm{y})-I_{\eta}\lvert^{2}\,{\rm d}\bm{y}\Big{)}^{\frac{1}{2}}\Big{\rvert}$
$=\Big{\lvert}\Big{(}\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}\lvert\eta(\bm{x})-\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}\eta(\bm{x})\,{\rm d}\bm{x}\lvert^{2}\,{\rm d}\bm{x}\Big{)}^{\frac{1}{2}}\Big{\rvert}$
$\leq\Big{\lvert}K_{\varepsilon}\lVert\nabla\eta(\bm{x})\lVert_{L^{2}(C_{\varepsilon})}\Big{\rvert}$
@@
$\square$
Proof of Remark 2.2. For any $\varepsilon>0$, let $\delta^{\prime}>0$ and $\delta>0$ such that $0<\lvert\bm{k}\rvert<\delta^{\prime}$ and $0<\lvert-\bm{k}^{2}-2\bm{k_{j}}\cdot\bm{k}\rvert<\delta$. Then,
$\Big{\lvert}\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}(\nabla^{2}+2\imath\bm{k_{j}}\cdot\nabla)\widehat{\eta_{j}}e^{\imath\bm{k}\cdot\frac{\bm{x}}{\varepsilon}}\,{\rm d}\bm{x}\Big{\rvert}$ $=\Big{\lvert}\int_{C_{1}}(\nabla^{2}+2\imath\bm{k_{j}}\cdot\nabla)\widehat{\eta_{j}}e^{\imath\bm{k}\cdot\bm{y}}\,{\rm d}\bm{y}\Big{\rvert}$
@@
$\leq\Big{\lvert}\Big{(}\int_{C_{1}}\lvert(-\bm{k}^{2}-2\bm{k_{j}}\cdot\bm{k})\widehat{\eta_{j}}\lvert^{2}\,{\rm d}\bm{y}\Big{)}^{\frac{1}{2}}\Big{(}\int_{C_{1}}\lvert e^{\imath\bm{k}\cdot\bm{y}}\lvert^{2}\,{\rm d}\bm{y}\Big{)}^{\frac{1}{2}}\Big{\rvert}$
$=\Big{\lvert}\Big{(}\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}\lvert(-\bm{k}^{2}-2\bm{k_{j}}\cdot\bm{k})\widehat{\eta_{j}}\lvert^{2}\,{\rm d}\bm{x}\Big{)}^{\frac{1}{2}}\Big{\rvert}$
$\leq\Big{\lvert}K_{\varepsilon}(-\bm{k}^{2}-2\bm{k_{j}}\cdot\bm{k})\lVert\nabla\widehat{\eta_{j}}\lVert_{L^{2}(C_{\varepsilon})}\Big{\rvert}$
$<\Big{\lvert}K_{\varepsilon}\delta\lVert\nabla\widehat{\eta_{j}}\lVert_{L^{2}(C_{\varepsilon})}\Big{\rvert}$
@@
$\square$
Appendix C Finite Element Methods for Solving (2.13)
Although we use the Fourier pseudospectral method to solve the APFC model in this work, we include a brief overview of the commonly used finite element method for completeness. This overview, based on previous studies [30, 24, 53, 54], focuses on spatial discretization and adaptivity techniques. For the discretization of the equation (2.13), a conforming triangulation $\mathcal{T}_{h}$ of the computational domain $\Omega$ is considered, usually with simplex elements $S\in\mathcal{T}_{h}$ of characteristic size $h$. In the simulation of APFC, linear elements are adopted. Considering a ($\mathbb{P}_{1}$) FEM space, we define
$\mathcal{V}_{h}^{1}=\{v\in\mathbb{C}(\Omega,\mathbb{R}):v|_{S}\in\mathbb{P}_{1}(S,\mathbb{R}),\,S\in\mathcal{T}_{h}\}.$
A function $y\in\mathcal{V}_{h}^{1}$ can be expressed in terms of a basis expansion: $y=\sum_{i}Y_{i}e_{i}$, with real coefficients $Y_{i}$ and basis $\{e_{i}\}$ of $\mathcal{V}_{h}^{1}$. A key feature of the FEM approach is splitting equation (2.13) into two second-order equations for $\partial\eta_{j}/\partial t$ and $\rho_{j}=\varsigma_{j}\eta_{j}$:
$\frac{\partial\eta_{j}}{\partial t}$ $=-|\bm{k}_{j}|^{2}\left(B_{x}\varsigma_{j}\rho_{j}+B_{0}\eta_{j}+3v\left(A^{2}-|\eta_{j}|^{2}\right)\eta_{j}+\frac{\partial f_{s}}{\partial\eta_{j}^{*}}\right),$ (C.1)
@@
Considering that the amplitude functions are complex-valued, we may use complex coefficients with real basis functions. The functions are split into real and imaginary parts and divided accordingly for numerical solution. The following array of functions is considered: $\alpha=\left[\mathrm{Re}(\eta_{1}),\mathrm{Im}(\eta_{1}),\ldots,\mathrm{Re}(\eta_{p}),\mathrm{Im}(\eta_{p})\right],\quad p=1,\ldots,2N$. The problem to solve then reads: for $t\in[0,T]$, find $\eta_{j}=\alpha_{k}+i\alpha_{k+1}$ and $\varsigma_{j}\eta_{j}=\zeta_{k}+i\zeta_{k+1}$, $k=2j-1$, with $\alpha_{k},\alpha_{k+1},\zeta_{k},\zeta_{k+1}\in\mathcal{V}_{h}^{1}$, such that
$\frac{\partial\alpha_{k}}{\partial t}$ $=-|\bm{k_{j}}|^{2}\bigg{[}B_{0}\alpha_{k}+B_{x}\left(\nabla^{2}\zeta_{k}-2\bm{k_{j}}\cdot\nabla\zeta_{k+1}\right)$ (C.2)
$\quad+3v\left(A^{2}-|\eta_{j}|^{2}\right)\alpha_{k}+\mathrm{Re}\left(\frac{\delta f_{s}}{\delta\eta_{j}^{*}}\right)\bigg{]},$
$\frac{\partial\alpha_{k+1}}{\partial t}$ $=-|\bm{k_{j}}|^{2}\bigg{[}B_{0}\alpha_{k+1}+B_{x}\left(\nabla^{2}\zeta_{k+1}+2\bm{k_{j}}\cdot\nabla\zeta_{k}\right)$
$\quad+3v\left(A^{2}-|\eta_{j}|^{2}\right)\alpha_{k+1}+\mathrm{Im}\left(\frac{\delta f_{s}}{\delta\eta_{j}^{*}}\right)\bigg{]},$
@@
The time derivatives are approximated by $\partial\alpha_{k}/\partial t=(a_{k}^{n+1}-a_{k}^{n})/\Delta t_{n}$ and $\partial\zeta_{k}/\partial t=(\zeta_{k}^{n+1}-\zeta_{k}^{n})/\Delta t_{n}$, with $\Delta t_{n}=t_{n+1}-t_{n}$ being the time step, and $n\in\mathbb{N}_{0}$ the index labeling time steps. The time discretization is achieved through a semi-implicit scheme, where the linear terms in (C.2) are evaluated implicitly and the nonlinear terms explicitly.
Notice that $N$ coupled systems must be solved concurrently, with $N$ being the number of independent amplitudes according to the considered lattice symmetry. The adopted semi-implicit integration scheme in matrix form reads:
@@
$\mathbf{L}$ $=\begin{bmatrix}-\nabla^{2}&\mathcal{A}&1&0\\ -\mathcal{A}&-\nabla^{2}&0&1\\ G_{1}(\{\alpha_{i}^{(n)}\})&0&\mathcal{K}\nabla^{2}&-\mathcal{K}\mathcal{A}\\ 0&G_{2}(\{\alpha_{i}^{(n)}\})&\mathcal{K}\mathcal{A}&\mathcal{K}\nabla^{2}\end{bmatrix},$ (C.3)
$\mathbf{x}$ $=\begin{bmatrix}\alpha_{k}^{(n+1)}\\ \alpha_{k+1}^{(n+1)}\\ \zeta_{k}^{(n+1)}\\ \zeta_{k+1}^{(n+1)}\end{bmatrix},\quad\mathbf{R}=\begin{bmatrix}0\\ 0\\ H_{1}(\{\alpha_{i}^{(n)}\})\\ H_{2}(\{\alpha_{i}^{(n)}\})\end{bmatrix},$ (C.4)
where $\mathcal{A}=2\bm{k_{j}}\cdot\nabla$ and $\mathcal{K}=|\bm{k_{j}}|^{2}B_{x}$. The functions evaluated explicitly at time $t_{n}$ are given by
$G_{1}(\{\alpha_{i}\})$ $=\frac{1}{\tau_{n}}+|\bm{k_{j}}|^{2}B_{0}+3v|\bm{k_{j}}|^{2}\left(A^{2}+\alpha_{k}^{2}-\alpha_{k+1}^{2}\right),$ (C.5)
$G_{2}(\{\alpha_{i}\})$ $=\frac{1}{\tau_{n}}+|\bm{k_{j}}|^{2}B_{0}+3v|\bm{k_{j}}|^{2}\left(A^{2}+\alpha_{k+1}^{2}-\alpha_{k}^{2}\right),$
$H_{1}(\{\alpha_{i}\})$ $=\left[\frac{1}{\tau_{n}}+6|\bm{k_{j}}|^{2}v\alpha_{k}^{2}\right]\alpha_{k}-|\bm{k_{j}}|^{2}\mathrm{Re}\left(\frac{\delta f_{s}}{\delta\eta_{j}^{*}}\right),$
$H_{2}(\{\alpha_{i}\})$ $=\left[\frac{1}{\tau_{n}}+6|\bm{k_{j}}|^{2}v\alpha_{k+1}^{2}\right]\alpha_{k+1}-|\bm{k_{j}}|^{2}\mathrm{Im}\left(\frac{\delta f_{s}}{\delta\eta_{j}^{*}}\right).$
Furthermore, spatial adaptivity can be adopted to improve the efficiency of numerical simulation. There are many methods to choose from, such as those based on error estimates or indicators. Due to the presence of defects, the amplitude functions may oscillate differently. Based on this property, numerical methods have been developed that focus on the oscillation of the complex amplitudes. We find the region where the quantity $\lVert\nabla A^{2}\rVert_{F}$ exceeds an arbitrary threshold, where $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm. The method then considers a refinement of the computational grid to ensure proper characterization of all the $\eta_{j}$ functions.

### v3

@@ B.3 Proofs of the Remarks in Section 2
Proof of Remark 1.
$\quad\Big\lvert\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}\eta(\bm{x})e^{\imath\bm{k}_{j}\cdot\frac{\bm{x}}{\varepsilon}}\,{\rm d}\bm{x}-\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}I_{\eta}e^{\imath\bm{k}_{j}\cdot\frac{\bm{x}}{\varepsilon}}\,{\rm d}\bm{x}\Big\rvert$
$=\Big\lvert\int_{C_{1}}(\eta(\varepsilon\bm{y})-I_{\eta})e^{\imath\bm{k}_{j}\cdot\bm{y}}\,{\rm d}\bm{y}\Big\rvert$
$\leq\Big\lvert\Big(\int_{C_{1}}\lvert\eta(\varepsilon\bm{y})-I_{\eta}\rvert^{2}\,{\rm d}\bm{y}\Big)^{\frac{1}{2}}\Big(\int_{C_{1}}\lvert e^{\imath\bm{k}_{j}\cdot\bm{y}}\rvert^{2}\,{\rm d}\bm{y}\Big)^{\frac{1}{2}}\Big\rvert$
$=\Big\lvert\Big(\int_{C_{1}}\big\lvert\eta(\varepsilon\bm{y})-I_{\eta}\big\rvert^{2}\,{\rm d}\bm{y}\Big)^{\frac{1}{2}}\Big\rvert$
$=\Big\lvert\Big(\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}\lvert\eta(\bm{x})-\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}\eta(\bm{x})\,{\rm d}\bm{x}\rvert^{2}\,{\rm d}\bm{x}\Big)^{\frac{1}{2}}\Big\rvert$
$\leq\Big\lvert K_{\varepsilon}\lVert\nabla\eta(\bm{x})\rVert_{L^{2}(C_{\varepsilon})}\Big\rvert$
@@
Proof of Remark 2. For any $\varepsilon>0$, let $\delta^{\prime}>0$ and $\delta>0$ such that $0<\lvert\bm{k}\rvert<\delta^{\prime}$ and $0<\lvert-\bm{k}^{2}-2\bm{k}_{j}\cdot\bm{k}\rvert<\delta$. Then,
$\quad\Big\lvert\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}(\nabla^{2}+2\imath\bm{k}_{j}\cdot\nabla)\widehat{\eta}_{j}e^{\imath\bm{k}\cdot\frac{\bm{x}}{\varepsilon}}\,{\rm d}\bm{x}\Big\rvert$
$=\Big\lvert\int_{C_{1}}(\nabla^{2}+2\imath\bm{k}_{j}\cdot\nabla)\widehat{\eta}_{j}e^{\imath\bm{k}\cdot\bm{y}}\,{\rm d}\bm{y}\Big\rvert$
@@
$\leq\Big\lvert\Big(\int_{C_{1}}\lvert(-\bm{k}^{2}-2\bm{k}_{j}\cdot\bm{k})\widehat{\eta}_{j}\rvert^{2}\,{\rm d}\bm{y}\Big)^{\frac{1}{2}}\Big(\int_{C_{1}}\lvert e^{\imath\bm{k}\cdot\bm{y}}\rvert^{2}\,{\rm d}\bm{y}\Big)^{\frac{1}{2}}\Big\rvert$
$=\Big\lvert\Big(\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}\lvert(-\bm{k}^{2}-2\bm{k}_{j}\cdot\bm{k})\widehat{\eta}_{j}\rvert^{2}\,{\rm d}\bm{x}\Big)^{\frac{1}{2}}\Big\rvert$
$\leq\Big\lvert K_{\varepsilon}(-\bm{k}^{2}-2\bm{k}_{j}\cdot\bm{k})\lVert\nabla\widehat{\eta}_{j}\rVert_{L^{2}(C_{\varepsilon})}\Big\rvert$
$<\Big\lvert K_{\varepsilon}\delta\lVert\nabla\widehat{\eta}_{j}\rVert_{L^{2}(C_{\varepsilon})}\Big\rvert$
@@
Appendix C Finite Element Methods for Solving (13)
Although we use the Fourier pseudospectral method to solve the APFC model in this work, we include a brief overview of the commonly used finite element method for completeness. This overview, based on previous studies 2017Controlling ; wang2021posteriori ; ortner2023framework , focuses on spatial discretization and adaptivity techniques. For the discretization of the equation (13), a conforming triangulation $\mathcal{T}_{h}$ of the computational domain $\Omega$ is considered, usually with simplex elements $S\in\mathcal{T}_{h}$ of characteristic size $h$. In the simulation of APFC, linear elements are adopted. Considering a ($\mathbb{P}_{1}$) FEM space, we define
$\mathcal{V}_{h}^{1}=\{v\in C(\Omega,\mathbb{R}):v\lvert_{S}\in\mathbb{P}_{1}(S,\mathbb{R}),\,S\in\mathcal{T}_{h}\}.$
A function $y\in\mathcal{V}_{h}^{1}$ can be expressed in terms of a basis expansion: $y=\sum_{i}Y_{i}e_{i}$, with real coefficients $Y_{i}$ and basis $\{e_{i}\}$ of $\mathcal{V}_{h}^{1}$. A key feature of the FEM approach is splitting equation (13) into two second-order equations for $\partial\eta_{j}/\partial t$ and $\rho_{j}=\varsigma_{j}\eta_{j}$:
$\frac{\partial\eta_{j}}{\partial t}$ $=-\lvert\bm{k}_{j}\rvert^{2}\left(B_{x}\varsigma_{j}\rho_{j}+B_{0}\eta_{j}+3v\left(A^{2}-\lvert\eta_{j}\rvert^{2}\right)\eta_{j}+\frac{\delta f_{s}}{\delta\eta_{j}^{*}}\right),$ (59)
@@ Appendix C Finite Element Methods for Solving (13)
Considering that the amplitude functions are complex-valued, we may use complex coefficients with real basis functions. The functions are split into real and imaginary parts and divided accordingly for numerical solution. The following array of functions is considered: $\alpha=\left[\mathrm{Re}(\eta_{1}),\mathrm{Im}(\eta_{1}),\ldots,\mathrm{Re}(\eta_{M}),\mathrm{Im}(\eta_{M})\right]$, indexed by $p=1,\ldots,2M$. The problem to solve then reads: for $t\in[0,T]$, find $\eta_{j}=\alpha_{k}+i\alpha_{k+1}$ and $\varsigma_{j}\eta_{j}=\zeta_{k}+i\zeta_{k+1}$, k=2j1k=2j-1, with $\alpha_{k},\alpha_{k+1},\zeta_{k},\zeta_{k+1}\in\mathcal{V}_{h}^{1}$, such that
$\frac{\partial\alpha_{k}}{\partial t}$ $=-\lvert\bm{k}_{j}\rvert^{2}\bigg[B_{0}\alpha_{k}+B_{x}\left(\nabla^{2}\zeta_{k}-2\bm{k}_{j}\cdot\nabla\zeta_{k+1}\right)$ (60)
$\quad+3v\left(A^{2}-\lvert\eta_{j}\rvert^{2}\right)\alpha_{k}+\mathrm{Re}\left(\frac{\delta f_{s}}{\delta\eta_{j}^{*}}\right)\bigg],$
$\frac{\partial\alpha_{k+1}}{\partial t}$ $=-\lvert\bm{k}_{j}\rvert^{2}\bigg[B_{0}\alpha_{k+1}+B_{x}\left(\nabla^{2}\zeta_{k+1}+2\bm{k}_{j}\cdot\nabla\zeta_{k}\right)$
$\quad+3v\left(A^{2}-\lvert\eta_{j}\rvert^{2}\right)\alpha_{k+1}+\mathrm{Im}\left(\frac{\delta f_{s}}{\delta\eta_{j}^{*}}\right)\bigg],$
@@
The time derivatives are approximated by $\partial\alpha_{k}/\partial t=(\alpha_{k}^{n+1}-\alpha_{k}^{n})/\Delta t_{n}$ and $\partial\zeta_{k}/\partial t=(\zeta_{k}^{n+1}-\zeta_{k}^{n})/\Delta t_{n}$, with $\Delta t_{n}=t_{n+1}-t_{n}$ being the time step, and $n\in\mathbb{N}_{0}$ the index labeling time steps. The time discretization is achieved through a semi-implicit scheme, where the linear terms in (60) are evaluated implicitly and the nonlinear terms explicitly.
Notice that $M$ coupled systems must be solved concurrently, with $M$ being the number of independent amplitudes according to the considered lattice symmetry. The adopted semi-implicit integration scheme in matrix form reads:
@@
$\mathbf{L}$ $=\begin{bmatrix}-\nabla^{2}&\mathcal{A}&1&0\\ -\mathcal{A}&-\nabla^{2}&0&1\\ G_{1}(\{\alpha_{i}^{(n)}\})&0&\mathcal{K}\nabla^{2}&-\mathcal{K}\mathcal{A}\\ 0&G_{2}(\{\alpha_{i}^{(n)}\})&\mathcal{K}\mathcal{A}&\mathcal{K}\nabla^{2}\end{bmatrix}$, (61)
$\mathbf{x}$ $=\begin{bmatrix}\alpha_{k}^{(n+1)}\\ \alpha_{k+1}^{(n+1)}\\ \zeta_{k}^{(n+1)}\\ \zeta_{k+1}^{(n+1)}\end{bmatrix},\quad\mathbf{R}=\begin{bmatrix}0\\ 0\\ H_{1}(\{\alpha_{i}^{(n)}\})\\ H_{2}(\{\alpha_{i}^{(n)}\})\end{bmatrix}$, (62)
where $\mathcal{A}=2\bm{k}_{j}\cdot\nabla$ and $\mathcal{K}=\lvert\bm{k}_{j}\rvert^{2}B_{x}$. The functions evaluated explicitly at time $t_{n}$ are given by
$G_{1}(\{\alpha_{i}\})$ $=\frac{1}{\Delta t_{n}}+\lvert\bm{k}_{j}\rvert^{2}B_{0}+3v\lvert\bm{k}_{j}\rvert^{2}\left(A^{2}+\alpha_{k}^{2}-\alpha_{k+1}^{2}\right),$ (63)
$G_{2}(\{\alpha_{i}\})$ $=\frac{1}{\Delta t_{n}}+\lvert\bm{k}_{j}\rvert^{2}B_{0}+3v\lvert\bm{k}_{j}\rvert^{2}\left(A^{2}+\alpha_{k+1}^{2}-\alpha_{k}^{2}\right),$
$H_{1}(\{\alpha_{i}\})$ $=\left[\frac{1}{\Delta t_{n}}+6\lvert\bm{k}_{j}\rvert^{2}v\alpha_{k}^{2}\right]\alpha_{k}-\lvert\bm{k}_{j}\rvert^{2}\mathrm{Re}\left(\frac{\delta f_{s}}{\delta\eta_{j}^{*}}\right),$
$H_{2}(\{\alpha_{i}\})$ $=\left[\frac{1}{\Delta t_{n}}+6\lvert\bm{k}_{j}\rvert^{2}v\alpha_{k+1}^{2}\right]\alpha_{k+1}-\lvert\bm{k}_{j}\rvert^{2}\mathrm{Im}\left(\frac{\delta f_{s}}{\delta\eta_{j}^{*}}\right).$
Furthermore, spatial adaptivity can be adopted to improve the efficiency of numerical simulations. There are many methods to choose from, such as those based on error estimates or indicators. Due to the presence of defects, the amplitude functions may oscillate differently. Based on this property, numerical methods have been developed that focus on the oscillation of the complex amplitudes. We find the region where the quantity $\lVert\nabla A^{2}\rVert_{F}$ exceeds an arbitrary threshold, where $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm. The method then considers a refinement of the computational grid to ensure proper characterization of all the $\eta_{j}$ functions.
Acknowledgements.
We are deeply grateful to Prof. Siran Li (Shanghai Jiao Tong University) for insightful discussions and valuable suggestions that greatly improved this paper, particularly for drawing our attention to the analysis of the analyticity of the APFC model and for pointing us to several helpful references on this topic.
Funding
X. Wei, Y. Wang and L. Zhang are partially supported by the National Natural Science Foundation of China (No. 12271360). K. Jiang is partially supported by the National Key Research and Development Program of China (No. 2023YFA1008802), the National Natural Science Foundation of China (No. 12171412), the Science and Technology Innovation Program of Hunan Province (No. 2024RC1052) and the Innovative Research Group Project of Natural Science Foundation of Hunan Province of China (No. 2024JJ1008). Y. Wang is also partially supported by the Development Postdoctoral Scholarship for Outstanding Doctoral Graduates from Shanghai Jiao Tong University. L. Zhang is also partially supported by the Shanghai Municipal Science and Technology Project (No. 23JC1402300) and the Fundamental Research Funds for the Central Universities.
Data Availability
Data sharing is not applicable to this paper as no datasets were generated or analyzed during the current study.
Conflict of interest
The authors declare that they have no conflict of interest.